# Generalized Short Path Algorithms: Towards Super-Quadratic Speedup over Markov Chain Search for Combinatorial Optimization

### v1

@@ 5.2.3 Sherrington-Kirkpatrick Model
Consider the (possibly diluted) Sherrington-Kirkpatrick Hamiltonian on a graph $G(\mathcal{N},\mathcal{E})$,
$H(x)=\frac{1}{\sqrt{n}}\sum_{(i,j)\in\mathcal{E}}^{n}g_{ij}x_{i}x_{j},$ (43)
@@
using Glauber dynamics.
@@
Let $P$ be a Glauber dynamics chain for SK model on a graph $G(\mathcal{N},\mathcal{E})$ with constant maximum degree, at inverse temperature $\beta<\beta_{c}$. Then, the log-Sobolev constant satisfies $\omega\geq\Omega(1/(n\log n))$.
@@
By [AJK+22, Theorem 12, part (a)], the modified Log Sobolev constant for Glauber dynamics is $\Omega(1/n)$ when $\beta<\beta_{c}$ (See the discussion in Page 11). On a graph with constant bounded degree, the transition probability of Glauber dynamics $\Theta(n^{-1})$. Hence, by Theorem 4.18, the log-Sobolev constant scales as $\Omega(1/(n\log(n)))$. ∎
@@
Let $P$ be a Glauber dynamics chain for the Sherrington-Kirkpatrick model on a regular graph $G(\mathcal{N},\mathcal{E})$ with a bounded maximum degree $d$ at inverse temperature parameter $\beta<\beta_{c}$ and stationary distribution $\pi_{\beta}$. Then there exists a short-path algorithm that finds the optimal solution of the Sherrington Kirkpatrick Hamiltonian with running time
@@
We first show that the tail bound holds for SK model. By using proposition 4 in [DPCB23], we know that the number of low energy states with energy smaller than $E^{\ast}(1-\eta)$ is smaller than $2^{\gamma n}$ where $\gamma$ is a constant. By assuming that $\log(1/\pi(E^{\ast}))=\Theta(n)$, we can conclude that the generalized tail bound holds as well. Note that if this assumption fails, then it means that there exists a sub-exponential solver for SK model. Finally, since $\lvert E^{\ast}\rvert=\Theta(n)$ and $\log(1/\pi(E^{\ast}))=\Theta(n)$ for the SK Model by using Lemma 5.18 and Theorem 4.9, we have $\gamma=\mathcal{O}(1)$. Thus, by Theorem 4.15, we have $b=\mathcal{O}(1/\log(n))$. Since $\Delta$ is constant by Lemma 5.17, the total runtime scales as $(\pi_{\beta}(E^{\ast}))^{-1\left(\frac{1}{2}-\frac{c}{\log(n)}\right)}$ due to Theorem 3.3. ∎
@@
Lemma 5.20.
Let $\pi$ be the Gibbs distribution corresponding to a cost function $H\colon\{0,1\}^{n}\mapsto\mathbb{R}$ at some positive inverse temperature $\beta>0$. We also assume that the cost function is concentrated away from its optimum, i.e., the number of states with cost greater than $(1-\eta)E^{\ast}$ is lower than $2^{-\gamma n}$.
@@
We first observe due to the assumption on concentration that the partition function $Z(\beta)\leq 2^{n}[(1-2^{-\gamma n})\exp(-\beta(1-\eta)E^{\ast})+2^{-\gamma n}\exp(-\beta E^{\ast})]$. As a consequence,
$\frac{\pi(E^{\ast})}{2^{-n}}$ $\geq\frac{\exp(-\beta E^{\ast})}{(1-2^{-\gamma n})\exp(-\beta(1-\eta)E^{\ast})+2^{-\gamma n}\exp(-\beta E^{\ast})}$ (44)
$\geq\frac{1}{0.5\exp(-\beta\eta|E^{\ast}|)+2^{-\gamma n}}$ (45)
$=\Omega\left(\max\left(2^{\gamma n},\exp(\beta\eta|E^{\ast}|)\right)\right).$ (46)
@@ 6 Numerical Results
First, we want to identify what values of $b$ are practically appropriate. In [DPCB23], the authors numerically show that the original short-path algorithm works well for the 3-spin problem for $b$ up to around 0.8, which is much larger than the theoretical bound of $b\leq 1.02\times 10^{-4}$. Here, we show that a similar observation can be made for the constrained and penalized cases. For (MaxCut-Hamming), Figure 1 A shows the quartiles of $b$ values that minimize the effective runtime (Equation 9) of the algorithm. Note that the $b$ values are hard-capped at 1.25 and may be higher. We see that as $n$ increases, the optimal $b$ converges to a range approximately between 0.8 and 1.2. However, when choosing the value of $b$ that needs to work for all instances, we want a conservative value that avoids encountering the possibly superexponentially small spectral gap. For this purpose, we identify the value of $b$ at which the phase transition occurs. We numerically characterize the phase transition point by the overlap of |$\ket{\psi_{b}}$ with the initial state dropping below 0.99. An example of the overlap with varying $b$ is shown in Figure 1 C. Empirically, we observe that the phase transition $b$ converges to around $0.78$ as $n$ increases. In Figure 1 B, we plot the quartiles of $b$ values that minimize the spectral gap. For most instances, the spectral gap is minimized when $b$ is greater than the phase transition value $\approx 0.78$. Therefore, we expect a ubiquitous value of phase transition to work for a (MaxCut-Hamming) instance with high probability.
@@
In Figure 3, we show the empirical worst-case scaling for all three problems with different choices of $b$. A proper selection of $b$ yields a super-Grover speedup across all examined problems. Conversely, when $b$ is excessively high, the algorithm may encounter a small spectral gap in the worst case. To demonstrate this, we use the runtime (Equation 9, which includes the inverse of the spectral gap term) as the metric and observe that the quality of the fitting degrades. Our numerics lead to two interesting conceptual observations: firstly, as observed also by [DPCB23] the optimal choices of $b$ are well beyond what is predicted by the theoretical analysis. Secondly, in the case of (MaxCut-Hamming) we numerically observe an advantage over quadratic speedup that does not decay with $n$ which is beyond the current theoretical analysis and indicates that the runtime of the long jump can possibly be characterized through weaker conditions than $\Delta_{p}$ smoothness. Finally, we confirm in our setting that is indeed reasonable to make the choice of $b$ by choosing the largest such value that allows for large overlap with the ground state. If the value of this critical $b$ asymptotes quickly as a function of $n$ this suggests a numerical mechanism for the development of efficient short-path algorithms.
Figure 3: Empirical fitting of the inverse overlap of the ground state $\lvert\langle\psi_{b}|z\rangle\rvert^{-1}$ and the runtime 9 of (MaxCut-Hamming), (MaxBisection), and MIS with difference choices of $b$. For the left column (MaxCut-Hamming), we use data with $n$ ranging from 10 to 30. Top: $b=0.8$ the fitted exponent is 0.392 with 95% confidence interval $[0.383,0.402]$, indicating there could exist $b$ that is better than the phase transition one in Figure 2; Bottom: $b=1$ the fitted exponent is 0.348 with 95% confidence interval $[0.271,0.426]$. For the middle column (MaxBisection), we use data with $n$ ranging from 16 to 22. Top: $b=0.7$ the fitted exponent is 0.444 with 95% confidence interval $[0.436,0.452]$; Bottom: $b=1$ the fitted exponent is 0.873 with 95% confidence interval $[0.842,0.905]$. For the right column (MIS), we use data with $n$ ranging from 10 to 21. Top: $b=0.6$ the fitted exponent is 0.400 with 95% confidence interval $[0.386,0.415]$; Bottom: $b=0.8$ the fitted exponent is 0.392 with 95% confidence interval $[0.122,0.663]$.
@@ Appendix A Technical Details for Generalized Short Path Framework
Lemma A.1 ($\theta$ short-path $\implies$ $\theta$ Spectral Gap Bound, adapted from Proposition 5 of [DPCB23]).
If $H_{b}$ satisfies the $\theta$ short-path condition, then the spectral gap of $H_{b}$ is at least $\theta$, i.e., all excited states have energy at least $-1+\theta$.
@@
Recall by construction, the ground state energy of $H_{b}$ is at most $-1$. Note that $\theta\leq 1$. Suppose, in order to arrive at a contradiction, that there are at least two orthogonal eigenstates $|\psi_{1}\rangle$ and $|\psi_{2}\rangle$, with energy strictly below $-1+\theta$. Since the short-path condition is satisfied, at least one of $|\psi_{1}\rangle$ or $|\psi_{2}\rangle$ has nonzero overlap with $|\pi\rangle$. Without loss of generality, assume this is the case for $|\psi_{2}\rangle$.
@@
Lemma A.2 (Runtime bound by approximate projector).
Let $\mathcal{M}=(\mathcal{X},P,\pi)$ be a reversible Markov chain with $|\mathcal{X}|=V$. Define
$\mathcal{P}_{\ell}:=\left(\frac{H_{b}}{\lvert E_{b}\rvert}\right)^{\ell},$
and let $\omega$ be the log-Sobolev constant of $D(P)$. For either all even or all odd $\ell$ we have
$\left\langle\sqrt{\pi}\middle|\mathcal{P}_{\ell}\middle|z\right\rangle-V\left(1-\frac{\omega}{2}\right)^{\ell}<\langle\sqrt{\pi}|\psi_{b}\rangle\left\langle\psi_{b}\middle|z\right\rangle.$
@@
Note that $E_{b}\leq-1$. By Lemma 4.12 all excited states have energy at least $-1+\frac{\omega}{2}$.
Since $\omega$ is the log-Sobolev constant of $D(P)$ (lower bounding its spectral gap $\delta$ as defined in (5)) and $g_{\eta}(H)$ is negative definite, at most two eigenvectors of $H_{b}$ have have magnitude $>1-\frac{\omega}{2}$, the ground state and highest-energy state. Thus,
$\langle\sqrt{\pi}|\mathcal{P}_{\ell}|z\rangle$ $=\langle\sqrt{\pi}|\psi_{b}\rangle\langle\psi_{b}|z\rangle+\sum_{|E_{b}^{\prime}|>1-\frac{\omega}{2}}\left(\frac{E^{\prime}_{b}}{E_{b}}\right)^{\ell}\langle\sqrt{\pi}|\psi^{\prime}_{b}\rangle\langle\psi^{\prime}_{b}|z\rangle+\sum_{|E_{b}^{\prime}|\leq 1-\frac{\omega}{2}}\left(\frac{E^{\prime}_{b}}{E_{b}}\right)^{\ell}\langle\sqrt{\pi}|\psi^{\prime}_{b}\rangle\langle\psi^{\prime}_{b}|z\rangle$
$\leq\langle\sqrt{\pi}|\psi_{b}\rangle\langle\psi_{b}|z\rangle+\left(\frac{E^{\prime}_{b}}{E_{b}}\right)^{\ell}\langle\sqrt{\pi}|\psi^{\prime}_{b}\rangle\langle\psi^{\prime}_{b}|z\rangle+(V-2)(1-\frac{\omega}{2})^{\ell}.$
Supposing $\ell$ takes a value such that the middle term is non-positive, we obtain
$\langle\sqrt{\pi}|\mathcal{P}_{\ell}|z\rangle-V(1-\frac{\omega}{2})^{\ell}<\langle\sqrt{\pi}|\psi_{b}\rangle\langle\psi_{b}|z\rangle.$
Lemma A.3 (ground-state energy shift bound, adapted from Proposition 6 of [DPCB23]).
Suppose that $\gamma$-spectral density and $\frac{\omega}{2}$-short-path condition hold, then
$\lvert E_{b}\rvert<1+\frac{4(\pi(E^{\ast}))^{\gamma}}{\omega}.$ (47)
@@
In the proof of Proposition 6 of [DPCB23], the authors showed that
$E_{b}$ $=\langle\psi_{b}|H_{b}|\psi_{b}\rangle=-1-b\langle\sqrt{\pi}|G_{\eta}|\sqrt{\pi}\rangle-b^{2}\langle\sqrt{\pi}|G_{\eta}W_{b}G_{\eta}|\sqrt{\pi}\rangle,$ (48)
where
$W_{b}$ $:=(\Pi_{\perp}(H_{b}-E_{b})\Pi_{\perp})^{-1}=\Pi_{\perp}(H_{b}-E_{b})^{-1}\Pi_{\perp}$
$\Pi_{\perp}$ $:=I-|\sqrt{\pi}\rangle\langle\sqrt{\pi}|.$
Recall the short-path condition
$\textup{{GSE}}(\Pi_{\perp}H_{b}\Pi_{\perp})\geq-1+\frac{\omega}{2}.$
Thus, with $E_{b}\leq-1$,
$\textup{{GSE}}(\Pi H_{b}\Pi)\geq-1+\frac{\omega}{2}\geq E_{b}+\frac{\omega}{2}$
$\implies\textup{{GSE}}(\Pi_{\perp}(H_{b}-E_{b})\Pi_{\perp}))\geq\omega/2,$
so $\lVert W_{b}\rVert_{2}\leq\frac{2}{\omega}$ and as a consequence,
$\langle\sqrt{\pi}|G_{\eta}(\Pi(H_{b}-E_{b})\Pi)^{-1}G_{\eta}|\sqrt{\pi}\rangle\leq\frac{2}{\omega}\langle\sqrt{\pi}|G_{\eta}^{2}|\sqrt{\pi}\rangle.$
From here, applying Equation (48) yields
$E_{b}\geq-1-b\langle\sqrt{\pi}|G_{\eta}|\sqrt{\pi}\rangle-\frac{2b^{2}}{\omega}\langle\sqrt{\pi}|G_{\eta}^{2}|\sqrt{\pi}\rangle.$
By spectral density and definition of $G_{\eta}$:
$-\langle\sqrt{\pi}|G_{\eta}|\sqrt{\pi}\rangle$ $\leq(\pi^{\ast})^{\gamma}$
$\langle\sqrt{\pi}|G_{\eta}^{2}|\sqrt{\pi}\rangle$ $\leq(\pi^{\ast})^{\gamma}.$
Using $0\leq b\leq 1$, $\pi^{\ast}<1,\omega<1$, we have
$E_{b}\geq-1-(\pi^{\ast})^{\gamma}\left(b+b^{2}\frac{2}{\omega}\right)>-1-\frac{4(\pi^{\ast})^{\gamma}}{\omega}.$
Next we provide an alternative bound on the energy shift. At a high level, the goal is to show that the small ground state energy shift condition is satisfied for any point we can efficiently go to in the short jump. It removes the need for the short-path condition. Note, heuristically, under this large ground state overlap condition, the short-path condition reduces to the large gap condition:
$\textup{{GSE}}((I-|\sqrt{\pi}\rangle\langle\sqrt{\pi}|)H_{b}(I-|\sqrt{\pi}\rangle\langle\sqrt{\pi}|))-E_{b}\approx\textup{{GSE}}((I-|\psi_{b}\rangle\langle\psi_{b}|)H_{b}(I-|\psi_{b}\rangle\langle\psi_{b}|))-E_{b}=\text{Gap}(H_{b}).$
Lemma A.4 ($\Delta_{P}(\eta)$ stable $\iff\alpha_{P}$ subdepolarizing ).
Let $\mathcal{M}=(\mathcal{X},P,\pi)$ be a reversible Markov chain. The pair $(\mathcal{M},H)$ satisfies the $\Delta_{P}(\eta)$-stable if and only if it satisfies $\alpha_{P}$-subdepolarizing. They are related by the equation $\alpha_{P}=\frac{\Delta_{P}(\eta)}{\lvert E^{\ast}\rvert(1-\eta)}$.
@@
$\prod_{t=1}^{T}f\left(\frac{c_{t}H(x)}{E^{\ast}}\left(1+\frac{\Delta}{H(x)}\right)\right)\geq\prod_{t=1}^{T}f\left(\frac{c_{t}H(x)}{E^{\ast}}\left(1-\frac{\Delta}{(1-\eta)\lvert E^{*}\rvert}\right)\right).$
@@
$\sum_{y}P(x,y)g_{\eta}\left(\lvert E^{*}\rvert^{-1}H(y)\right)\leq g_{\eta}\left(\lvert E^{*}\rvert^{-1}(H(x)+\alpha_{P}(1-\eta)\lvert E^{*}\rvert)\right).$
Lemma A.5 (Upper bounds on $\Delta(\eta)$).
Let $\eta\in[0,1)$. If $H$ has $P$ pseudo-Lipschitz norm $\lVert H\rVert_{P}$, then
$\sqrt{\lVert H\rVert_{P}}\geq\Delta_{P}(\eta).$
Furthermore, if
@@
which may actually be easier to show, then we also have $\Delta_{P}(\eta)$ stability with $\Delta_{P}(\eta)\leq\tilde{\Delta}_{P}$, since by concavity of $h_{\eta}=g_{\eta}(\frac{x}{\lvert E^{*}\rvert})$ and Jensen’s inequality:
[…]

### v2

@@ 5.2.4 Sherrington-Kirkpatrick Model
Consider the Sherrington-Kirkpatrick Hamiltonian
$H(x)=\frac{1}{\sqrt{n}}\sum_{i<j\leq n}^{n}g_{ij}x_{i}x_{j},$ (40)
@@
using Glauber dynamics. The following provides a bound on $\Delta_{P}$ for Glauber over SK.
@@
Let $P$ be a Glauber dynamics chain for SK model on a graph $G(\mathcal{N},\mathcal{E})$ at inverse temperature $\beta<\frac{1}{4}$. Then, the log-Sobolev constant satisfies $\omega\geq\Omega(1/(n\log n))$.
@@
By [CE22, Corollary 51], the modified log-Sobolev constant for Glauber dynamics is $\Omega(1/n)$ when $\beta<\frac{1}{4}$. For possible transitions, the transition probability of Glauber dynamics is $\Omega(n^{-1})$. Hence, by Theorem 4.17, the log-Sobolev constant scales as $\Omega(1/(n\log n))$. ∎
@@
Let $P$ be a Glauber dynamics chain for the Sherrington-Kirkpatrick model at inverse temperature parameter $\beta<\frac{1}{4}$ and stationary distribution $\pi_{\beta}$. Then, there exists a short path algorithm that finds the optimal solution of the Sherrington Kirkpatrick Hamiltonian with running time
@@
We first show that the tail bound holds for SK model. By using proposition 4 in [DPCB23], we know that the number of low energy states with energy smaller than $E^{\ast}(1-\eta)$ is smaller than $2^{\gamma n}$ where $\gamma$ is a constant. By assuming that $\log(1/\pi(E^{\ast}))=\Theta(n)$, we can conclude that the generalized tail bound holds as well. Note that if this assumption fails, then it means that there exists a sub-exponential solver for SK model. Finally, since $\lvert E^{\ast}\rvert=\Theta(n)$ and $\log(1/\pi(E^{\ast}))=\Theta(n)$. Thus, by Lemma 5.34 and Theorem 4.14, we have $b=\mathcal{O}(1/\log(n))$. Since $\Delta$ is constant by Lemma 5.33, the total runtime scales as $(\pi_{\beta}(E^{\ast}))^{-1\left(\frac{1}{2}-\frac{c}{\log(n)}\right)}$ due to Theorem 4.2. ∎
@@
Proposition 5.36.
Let $\pi_{\beta}$ be the Gibbs distribution corresponding to the SK cost function $H\colon\{0,1\}^{n}\rightarrow\mathbb{R}$ at some positive inverse temperature $\beta>0$. Then the probability of observing the ground state under the Gibbs measure is at least $2^{-\kappa n}$. Where, $\kappa>-1$ for sufficiently large $n$.
@@
Let $Z(\beta)$ be the partition function associated with $\pi_{\beta}$. We know that $\frac{\log Z(\beta)}{n}$ has subGaussian tails around its mean for every fixed $\beta$ [Tal10, Equation 1.54]. A loose version of Guerra’s replica symmetric bound [Tal10, Theorem 1.3.7] gives that
$\frac{\mathbb{E}[\log Z(\beta)]}{n}\leq 1+\frac{\mathbb{E}[\ln\cosh(\beta z)]}{\ln(2)}.$
Hence with high probability if $\frac{\lvert E^{\ast}\rvert}{n}=:\lvert P^{\ast}\rvert\approx.763\dots+o_{n}(1)$
$\frac{\log(\pi(E^{\ast}))}{n}\geq\frac{\beta}{\ln(2)}\left(\lvert P^{\ast}\rvert-\frac{\mathbb{E}[\ln\cosh(\beta z)]}{\beta}\right)-1+\epsilon(n),$
which is >1>-1 for sufficiently large $n$. ∎
@@ 6 Numerical Results
[…]
@@
In Figure 3, we show the empirical worst-case scaling for all three problems with different choices of $b$. A proper selection of $b$ yields a super-Grover speedup across all examined problems. Conversely, when $b$ is excessively high, the algorithm may encounter a small spectral gap in the worst case. To demonstrate this, we use the runtime (Equation 9, which includes the inverse of the spectral gap term) as the metric and observe that the quality of the fitting degrades. Our numerics lead to two interesting conceptual observations: firstly, as observed also by [DPCB23] the optimal choices of $b$ are well beyond what is predicted by the theoretical analysis. Secondly, in the case of (MaxCut-Hamming) we numerically observe an advantage over quadratic speedup that does not decay with $n$ which is beyond the current theoretical analysis and indicates that the runtime of the long jump can possibly be characterized through weaker conditions than $\Delta_{P}$ stability. Finally, we confirm in our setting that is indeed reasonable to make the choice of $b$ by choosing the largest such value that allows for large overlap with the ground state. If the value of this critical $b$ asymptotes quickly as a function of $n$ this suggests a numerical mechanism for the development of efficient short path algorithms.
Figure 3: Empirical fitting of the inverse overlap of the ground state $\lvert\langle\psi_{b}|z\rangle\rvert^{-1}$ and the runtime 9 of (MaxCut-Hamming), (MaxBisection), and MIS with difference choices of $b$. For the left column (MaxCut-Hamming), we use data with $n$ ranging from 10 to 30. Top: $b=0.8$ the fitted exponent is 0.392 with 95% confidence interval $[0.383,0.402]$, indicating there could exist $b$ that is better than the phase transition one in Figure 2; Bottom: $b=1$ the fitted exponent is 0.348 with 95% confidence interval $[0.271,0.426]$. For the middle column (MaxBisection), we use data with $n$ ranging from 16 to $2$. Top: $b=0.7$ the fitted exponent is 0.444 with 95% confidence interval $[0.436,0.452]$; Bottom: $b=1$ the fitted exponent is 0.873 with 95% confidence interval $[0.842,0.905]$. For the right column (MIS), we use data with $n$ ranging from 10 to 21. Top: $b=0.6$ the fitted exponent is 0.400 with 95% confidence interval $[0.386,0.415]$; Bottom: $b=0.8$ the fitted exponent is 0.392 with 95% confidence interval $[0.122,0.663]$.
@@ Appendix A Technical Details for Generalized Short Path Framework
[…]
@@
Recall by construction, the ground state energy of $H_{b}$ is at most 1-1. Note that $\theta\leq 1$. Suppose, in order to arrive at a contradiction, that there are at least two orthogonal eigenstates $|\psi_{1}\rangle$ and $|\psi_{2}\rangle$, with energy strictly below $1+\theta$. Since the short path condition is satisfied, at least one of $|\psi_{1}\rangle$ or $|\psi_{2}\rangle$ has nonzero overlap with $|\pi\rangle$. Without loss of generality, assume this is the case for $|\psi_{2}\rangle$.
@@
[…]
@@
[…]
$|\sqrt{\pi}\rangle=\langle\sqrt{\pi}|{\psi_{b}}\rangle|\psi_{b}\rangle+\sqrt{1-\langle\sqrt{\pi}|{\psi_{b}}\rangle^{2}}|\psi_{b}^{\perp}\rangle,$
where by stoquasticity of $H_{b}$, $\langle\sqrt{\pi}|\psi_{b}\rangle\geq 0$.
The short path condition implies that
$\langle\psi_{b}^{\perp}|H_{b}|\psi_{b}^{\perp}\rangle\geq-1+\theta,$ (41)
and since $\langle\psi_{b}|H_{b}|\psi_{b}\rangle\leq-1$, we have
$\lvert\langle\psi_{b}^{\perp}|H_{b}|\psi_{b}^{\perp}\rangle-\langle\psi_{b}|H_{b}|\psi_{b}\rangle\rvert\geq\theta.$ (42)
Combining the above with $\lVert H_{b}\rVert_{2}\leq 2$ gives
$\theta^{2}\leq\lvert\langle\psi_{b}|H_{b}|\psi_{b}\rangle-\langle\psi_{b}^{\perp}|H_{b}|\psi_{b}^{\perp}\rangle\rvert^{2}$ $\leq 4\lVert|\psi_{b}\rangle-|\psi_{b}^{\perp}\rangle\rVert^{2}$
$=8(1-\langle\sqrt{\pi}|\psi_{b}^{\perp}\rangle)$
$\leq 8(\langle\sqrt{\pi}|\psi_{b}\rangle)^{2},$
so the result follows. ∎
[…]
@@
Define
$\mathcal{P}_{\ell}:=\left(\frac{H_{b}}{E_{b}}\right)^{\ell}.$
Since $P$ is aperiodic, we have $I\preceq-D(P)\prec I$. The negative definiteness of $g_{\eta}(H)$ implies the maximum eigenvalue of $H_{b}$ cannot exceed that of D(P)-D(P). Hence, since $\nu$ is a lower bound on either the log-Sobolev or Poincaré constant of $P$ (lower bounding its spectral gap as defined in (5)), we have that $H_{b}$ does not have any eigenstates with eigenvalue $1-\frac{\nu}{2}$. Additionally, the only eigenstate with energy below $1+\frac{\nu}{2}$ is the ground state $|\psi_{b}\rangle$, by assumption.
Thus,
$\langle\sqrt{\pi}|\mathcal{P}_{\ell}|z\rangle$ $=\langle\sqrt{\pi}|\psi_{b}\rangle\langle\psi_{b}|z\rangle+\sum_{|E_{b}^{\prime}|\leq 1-\frac{\nu}{2}}\left(\frac{E^{\prime}_{b}}{E_{b}}\right)^{\ell}\langle\sqrt{\pi}|\psi^{\prime}_{b}\rangle\langle\psi^{\prime}_{b}|z\rangle$
$\leq\langle\sqrt{\pi}|\psi_{b}\rangle\langle\psi_{b}|z\rangle+(V-1)(1-\frac{\nu}{2})^{\ell},$
as all $E_{b}^{\prime}<0$.
The above implies
$\lvert E_{b}\rvert^{-\ell}\cdot\langle\sqrt{\pi}|(-H_{b})^{\ell}|z\rangle-V(1-\frac{\nu}{2})^{\ell}<\langle\sqrt{\pi}|\psi_{b}\rangle\langle\psi_{b}|z\rangle<\lvert\langle\psi_{b}|z\rangle\rvert$.
We also have that
$\lvert E_{b}\rvert\langle\psi_{b}|\sqrt{\pi}\rangle$ $=\langle\psi_{b}|D(P)-bG_{\eta}|\sqrt{\pi}\rangle$
$=\langle\psi_{b}|\sqrt{\pi}\rangle-b\langle\psi_{b}|G_{\eta}|\sqrt{\pi}\rangle$
$\leq\langle\psi_{b}|\sqrt{\pi}\rangle+b\sqrt{\mathbb{P}_{\pi}(E\leq(1-\eta)E^{*})}.$
Since $\langle\psi_{b}|\sqrt{\pi}\rangle\geq 0$ by stoquasticity, we get
$\lvert E_{b}\rvert^{-\ell}\geq\exp\left(-\ell\frac{b\sqrt{\mathbb{P}_{\pi}(E\leq(1-\eta)E^{*})}}{\langle\psi_{b}|\sqrt{\pi}\rangle}\right).$
If $\ell$ satisfies the conditions in the Lemma statement, then there exists an integer giving the stated result. ∎
[…]
@@
[…]
@@
$\mathbb{E}_{y\underset{P}{\sim}x}[h_{\eta}\left(\lvert E^{*}\rvert^{-1}H(y)\right)]\leq h_{\eta}\left(\sum_{y}P(x,y)H(y)\right)\leq h_{\eta}(H(x)+\tilde{\Delta}_{P}).$